if there is a regular polygon and each angle is 170 degrees then how many sides does the polygon have?
Answer by Alan3354(69443)

if there is a regular polygon and each angle is 170 degrees then how many sides does the polygon have?
-------------------
Exterior angles = 180 - 170 = 10 degs
# of sides = 360/10 = 36